# MMLongBench: Benchmarking Long-Context Vision-Language Models Effectively and Thoroughly

### v2

@@ A.6 Image Resizing and Statistics
As we discussed in Section 3.3, the number of tokens per image is determined by the image size in our benchmark. In MM-NIAH, we find that many images from the OBELICS dataset are unnecessarily large (up to 8000$\times$6000 pixels) and are not text-rich. Then, we resize those images’ longer edge to 1024 pixels while preserving their aspect ratio.
We calculate the average number of images per example in all the datasets and input lengths in Table 3. From the table, we can find that our benchmark covers various text-to-image ratios. For example, VRAG tasks are text-centric and only contain one image per example, while ICL represents image-centric tasks with hundreds of images. MM-NIAH tasks are intermediate and feature both substantial text and multiple images.
@@ A.7 License
Table 3: Average number of images per example in all datasets and input lengths. The values in subscript denote the standard deviations. (T) and (I) represents the text and image needle in each task of MM-NIAH.
@@ Appendix B Full Model List
We list all 46 models [10, 43, 85, 34, 47, 29, 83, 30, 84, 3, 49, 50, 51, 44, 98, 99] we evaluated in Table 10. All 46 models have a pixel unshuffle operation to reduce the token counts of images. This is consistent with our token counting methods (Section 3.2). The only exception is Pixtral-12B, but we can resize its image (to 0.5$\times$ on each side) to reduce the image tokens. Thus, we can fit Pixtral-12B on our GPU server and avoid extremely long input sequences.
@@ B.1 LVLMs Beyond our Evaluation
Beyond the models in our evaluation, there are also a lot of excellent models, such as Llava-onevision [102], Llama3.2 [1], mPLUG-Owl3 [45]. However, we find that those models don’t have pixel unshuffle operations. While Pixtral-12B’s ViT can take images with dynamic resolution, the ViTs of these three models use dynamic tiling, such as $560\times 560$ tiles for Llama3.2-11B. Thus, unlike Pixtral-12B, we cannot reduce the number of image tokens by simply resizing the input image, since these models do not accept images smaller than the predefined tile size. As shown in Table 11, these models cannot efficiently process high-resolution images as a whole, generating a number of tokens that is at least ten times the pre-defined input length (64K or 128K). This causes extremely long input sequences, and we cannot fit them on our GPU server.
@@ Appendix C Experimental Setup
For all open-source models, we evaluate them on an 8$\times$×A100 (80GB) GPU server. We use the HuggingFace Transformers framework [104] to deploy models and generate outputs. Since all models are instruction-tuned, we apply the chat templates to all datasets. We load models in BF16 with FlashAttention2 [105] for faster inference. The largest open-source models tested in our work have 72B parameters. Our computational resources are limited to 8$\times$×A100 GPUs; thereby, we cannot evaluate models with over 100B parameters, such as Llama4 [4], at 128K tokens.
@@ Appendix D Additional Results
We provide more evaluation results in addition to Section 4.
@@ D.1 The Difficulty of Visual Haystack
We discussed the difficulty of Visual Haystack in Section 4.2. As shown in Figure 3, current LCVLMs achieve the performance only slightly higher than random guessing on VH-Multi. Here, in Figure 10, we present the performance of the selected models on both VH-Single and VH-Multi, providing a complete view. We find that models also perform poorly on VH-Single.
@@
We manually checked a number of examples from the Visual Haystack dataset and didn’t find any errors in the task labels. As shown in Figure 10, Gemini-2.5-Pro achieves an accuracy of 85.4 on VH-Single, and GPT-4o achieves an accuracy of 70.3 on VH-Multi. These results are higher than a random guess (50%), which demonstrates the correctness of our implementation.
@@ D.2 Correlation between NIAH and Various Downstream Tasks
We discussed the correlation between NIAH tasks and various downstream applications in Figure 4. Here, we provide a detailed version of the task correlation in Figure 11. For the three tasks in MM-NIAH, we also report the correlations on the subsets containing only text-needle or only image-needle examples. We can find that subsets of image-needle examples correlate less with various downstream tasks compared to text-needle examples, especially MM-NIAH-Count (I) and MM-NIAH-Reason (I)
We further show the performance of the selected models on the text-needle and image-needle subsets in Figure 12. We observe that models exhibit weak performance on MM-NIAH-Count (I) and MM-NIAH-Reason (I). This challenging nature leads to a low degree of separability between different models.
@@ D.3 Correlation between Datasets
We plot the correlation between all MMLongBench datasets and category averages in Figure 13. Generally, the datasets in each category strongly correlate with each other. The VH-Single and VH-Multi are exceptions, due to their high difficulty. Also, MM-NIAH-Count exhibits relatively weak correlations with MM-NIAH-Ret and MM-NIAH-Reason, suggesting that counting is a different skill from retrieving needles (key information) and subsequently reasoning over them.
@@ D.4 Performance of Claude
At the time of our evaluation, the Claude 3 family of models can take up to 100 images [^6] https://docs.anthropic.com/en/docs/build-with-claude/vision#basics-and-limits in a single request. However, a few datasets, such as Food101, VH-Single, or SlildeVQA, contain hundreds of images at input lengths of 64K and 128K tokens. As a result, it is impossible to process all images in a single pass through the model. For each input length, if one or more datasets within a category contain samples with more than 100 images, we exclude that category from evaluation at that input length. We provide the statistics about the average image number per example in each dataset at all five input lengths in Table 3 and Section A.6. When testing Claude-3.7-Sonnet, we mark those untestable cases as “N/A” in our results to distinguish them from genuine model failures (which receive a score of 0).
@@
D.5 Positional Embedding Extrapolation Experiments
@@
According to its technical reports [29], Qwen2.5-VL models are pre-trained with a context length of 32K tokens during the “Long-Context Pre-Training” stage. Meanwhile, its HuggingFace model card shows that we can use YaRN [37], with a scaling factor of 4, to extend its context length to 128K tokens [^7] https://huggingface.co/Qwen/Qwen2.5-VL-3B-Instruct#processing-long-texts. Note that YaRN is used in a zero-shot manner, as there is no continual training for YaRN on Qwen2.5-VL. In Figure 14, we test the performance of using YaRN. We have two observations: (1) Using YaRN may hurt the performance on shorter input lengths. For example, at 8K tokens, the DocVQA score of Qwen2.5-VL-32B decreases from 67.8% to 63.4%; (2) On average across the entire MMLongBench, YaRN only substantially improves the performance of Qwen2.5-VL-3B (from 21.2 to 30.2 at 128K). However, the SoTA performance from large models (i.e., 32B and 72B) only fluctuates slightly.
@@
Ge et al. [52] proposed a positional embedding extrapolation method called V2PE, where it assigns smaller positional increments to visual tokens than textual tokens. They further applied V2PE to InternVL2-2B and trained the model to enhance its performance on MM-NIAH-Ret (I). In this experiment, we evaluate the V2PE-256K checkpoints [^8] https://huggingface.co/OpenGVLab/V2PE/tree/main/V2PE-256K with the visual increments $\delta\in\{\frac{1}{16},\frac{1}{64},\frac{1}{256}\}$. As shown in Figure 15, we find that (1) V2PE is very sensitive to different visual increments. For example, when we use $\frac{1}{16}$ and $\frac{1}{64}$, the performance is even worse than InternVL2-2B, which leads to extra hyperparameter tuning; (2) The V2PE (256) shows extremely high performance on NIAH tasks, which can be attributed to the fact that it was trained on MM-NIAH. The sharp performance difference on NIAH tasks versus other categories suggests that V2PE is strongly overfitted to the MM-NIAH dataset.
D.6 Lost in the Middle
@@
Existing works found that text-pure LLMs often struggle to recall needles in the middle of the input sequence, named lost in the middle [57]. On our benchmark, we extend the previous analysis to vision-language tasks with input length up to 128K tokens. We place the needle at six different evenly spaced depths in the context (i.e., $[0,0.2,0.4,0.6,0.8,1.0]$) and evaluate the LCVLMs’ ability to retrieve it. In our study, the needle may be a gold passage, an image, or a key sentence. We show the results in Figure 16 for ViQuAE, Figure 17 for VH-Single, Figure 18 for MM-NIAH-Ret (T), Figure 19 for MM-NIAH-Ret (I), Figure 20 for MM-NIAH-Reason (I).
@@
We observe a similar lost-in-the-middle phenomenon in many LCVLMs on long-context vision-language tasks. For example, the InternVL3-14B in Figure 18 and Ovis2-34B in Figure 19 both exhibit much better performance when the needle is at depths 00 and $1.0$. Furthermore, as we extend the context to longer lengths (e.g., 128K tokens), we observe cases where the model tends to favor either the very beginning or the very end of the context, but not both simultaneously. For example, as shown in Figure 16, InternVL3-8B prefers the very beginning of the context (depth 00) at 128K tokens, whereas Qwen2.5-VL-72B favors the very end (depth $1.0$).
D.7 Error Analysis Details
@@
We conducted two error analyses in Section 4.4. We provide more details of those two analyses in this section. First, for MMLongBench-Doc, we used PyMuPDF  [^9] https://pymupdf.readthedocs.io to extract the plain text from PDF-formatted documents. For ViQuAE, the entity names are already provided in the dataset, since it is constructed based on TriviaQA. The text-pure LLMs we used, corresponding to Qwen2.5-VL models, are Qwen2.5-7B-Instruct and Qwen2.5-32B-Instruct, which are instruction-tuned versions.
D.8 Full Model Evaluation Results
@@
In Figure 21, we provide the results of all 46 models. We also plot the performance of all 46 models on each dataset in Figures 22, 23, 24 and 25.
D.9 Idefics2 Performance
@@
The Idefics2-8B and Idefics-8B-C only have a training context window of 2K tokens [49]. We find this leads to very poor long-context generalization. Also, the LLM used in Idefics2 is Mistral-7B-v0.1 [32], whose training length is only 8K tokens. From Figure 18, we observe that Idefics2 models perform well only when the needle depth is $1.0$ and the context is short (8K or 16K tokens). Additionally, we conduct a sanity check by removing all negative images and retaining only the needle images in Visual Haystack (i.e., one image for single-needle examples and two or three images for multi-needle examples). As shown in Table 12, we observe that both models achieve performance much higher than a random guess (50%), indicating the correctness of the implementation.
@@
Table 12: Sanity check of Idefics2-8B and Idefics2-8B-C. Here we use the Visual Haystack dataset. We remove all negative images and only retain needle images (i.e., one image for single-needle examples and two or three images for multi-needle examples).
@@ Appendix E Prompts and Data Examples
We list a few data examples with prompts in Figures 26, 27, 28, 29, 30 and 31. For NIAH tasks, we provide examples of both VH and MM-NIAH, as their input formats are very different.
@@ Appendix F Limitation
For limitation of evaluated models, while we already provide an extensive coverage of 46 frontier LCVLMs, there are still some models that we cannot cover due to token efficiency or integration challenges of codebases, as we discussed in Section B.1. We leave those works for future study. Meanwhile, the largest open-source models we evaluated are up to 72B in size (Qwen2-VL-72B and Qwen2.5-VL-72B). As we discussed in Appendix C, our computational resources are limited to 8$\times$×A100 (80G) GPUs; thereby it is hard to deploy and evaluate larger models with over 100B parameters at the input length of 128K tokens, such as Llama4 [4].
For evaluating summarization, we use a model-based metric (See Section A.4) that can provide much better alignment with human judgment than N-gram overlap metrics, such as ROUGE-L. However, we find using GPT-4o to provide the evaluation is expensive, which prevents the long-context community from conducting evaluations with hundreds or even thousands of models. Therefore, it is necessary to find an alternative evaluation method with a lower cost.
@@ Appendix G Broader Impact
The long-context ability of LVLMs has unlocked a large range of applications, including understanding documents with hundreds of pages and reasoning over dozens of web pages automatically. This ability can also help users to summarize a long document or revise a large-scale code repository. Meanwhile, there are a large number of instruction-following scenarios grounded in complex vision-language contexts, such as long-term dialogue with humans or dialogue-based navigation for robots. Looking ahead, our MMLongBench will serve as a standard evaluation for the whole community to benchmark new LCVLMs and to stimulate the development of models with more efficient vision-language token encodings, more robust position-extrapolation schemes, and improved OCR, multi-modal retrieval, and reasoning capabilities.
@@
Figure 28: Example of MM-NIAH-Ret dataset in NIAH category.

### v3

@@ A.6 Image Resizing and Statistics
[…]
We calculate the average number of images per example in all the datasets and input lengths in Table˜3. From the table, we can find that our benchmark covers various text-to-image ratios. For example, VRAG tasks are text-centric and only contain one image per example, while ICL represents image-centric tasks with hundreds of images. MM-NIAH tasks are intermediate and feature both substantial text and multiple images.
@@ A.7 License
Table 3: Average number of images per example in all datasets and input lengths. The values in subscript denote the standard deviations. (T) and (I) represents the text and image needle in each task of $M$-NIAH.
@@ Appendix B Full Model List
[…]
@@ B.1 LVLMs Beyond our Evaluation
Beyond the models in our evaluation, there are also a lot of excellent models, such as Llava-onevision [102], Llama3.2 [1], mPLUG-Owl3 [45]. However, we find that those models don’t have pixel unshuffle operations. While Pixtral-12B’s ViT can take images with dynamic resolution, the ViTs of these three models use dynamic tiling, such as $560\times 560$ tiles for Llama3.2-11B. Thus, unlike Pixtral-12B, we cannot reduce the number of image tokens by simply resizing the input image, since these models do not accept images smaller than the predefined tile size. As shown in Table˜11, these models cannot efficiently process high-resolution images as a whole, generating a number of tokens that is at least ten times the pre-defined input length (64K or 128K). This causes extremely long input sequences, and we cannot fit them on our GPU server.
@@ Appendix C Experimental Setup
[…]
@@ Appendix D Additional Results
We provide more evaluation results in addition to Section˜4.
@@ D.1 The Difficulty of Visual Haystack
We discussed the difficulty of Visual Haystack in Section˜4.2. As shown in Figure˜3, current LCVLMs achieve the performance only slightly higher than random guessing on VH-Multi. Here, in Figure˜12, we present the performance of the selected models on both VH-Single and VH-Multi, providing a complete view. We find that models also perform poorly on VH-Single.
@@
We manually checked a number of examples from the Visual Haystack dataset and didn’t find any errors in the task labels. As shown in Figure˜12, Gemini-2.5-Pro achieves an accuracy of 85.4 on VH-Single, and GPT-4o achieves an accuracy of 70.3 on VH-Multi. These results are higher than a random guess (50%), which demonstrates the correctness of our implementation.
@@ D.2 Correlation between NIAH and Various Downstream Tasks
We discussed the correlation between NIAH tasks and various downstream applications in Figure˜4. Here, we provide a detailed version of the task correlation in Figure˜13. For the three tasks in MM-NIAH, we also report the correlations on the subsets containing only text-needle or only image-needle examples. We can find that subsets of image-needle examples correlate less with various downstream tasks compared to text-needle examples, especially MM-NIAH-Count (I) and MM-NIAH-Reason (I)
We further show the performance of the selected models on the text-needle and image-needle subsets in Figure˜14. We observe that models exhibit weak performance on MM-NIAH-Count (I) and MM-NIAH-Reason (I). This challenging nature leads to a low degree of separability between different models.
@@ D.3 Correlation between Datasets
We plot the correlation between all MMLongBench datasets and category averages in Figure˜15. Generally, the datasets in each category strongly correlate with each other. The VH-Single and VH-Multi are exceptions, due to their high difficulty. Also, MM-NIAH-Count exhibits relatively weak correlations with MM-NIAH-Ret and MM-NIAH-Reason, suggesting that counting is a different skill from retrieving needles (key information) and subsequently reasoning over them.
@@ D.4 Performance of Claude
At the time of our evaluation, the Claude 3 family of models can take up to 100 images [^6] https://docs.anthropic.com/en/docs/build-with-claude/vision#basics-and-limits in a single request. However, a few datasets, such as Food101, VH-Single, or SlildeVQA, contain hundreds of images at input lengths of 64K and 128K tokens. As a result, it is impossible to process all images in a single pass through the model. For each input length, if one or more datasets within a category contain samples with more than 100 images, we exclude that category from evaluation at that input length. We provide the statistics about the average image number per example in each dataset at all five input lengths in Table˜3 and Section˜A.6. When testing Claude-3.7-Sonnet, we mark those untestable cases as “N/A” in our results to distinguish them from genuine model failures (which receive a score of 0).
D.5 More Models for Reasoning Comparison
In Section˜4.1, we observe that reasoning ability enhances multimodal long-context performance, as demonstrated by the Gemini model series. Here, we additionally evaluate MiMo-VL-7B-SFT (Reasoning) to further support this finding. As shown in Table˜12, we observe that the average performance of MiMo-VL-7B-SFT is much higher than Qwen2.5-VL-7B across all context lengths. At 8K context length, MiMo-VL-7B-SFT achieves 66.0 on average, substantially surpassing Qwen2.5-VL-7B’s 57.4. Meanwhile, MiMo-VL-7B-SFT consistently achieves higher performance on each task, except for summarization (this is quite common; see Section˜4.1, where different models exhibit varying strengths). This advantage also persists as the context length increases.
Table 12: Performance comparison of MiMo-VL-7B-SFT (Reasoning) and Qwen2.5-VL-7B (Non-Reasoning) with different context lengths.
Context Model VRAG NIAH ICL Summ DocVQA Avg.
8K MiMo-VL-7B-SFT 67.4 75.5 98.0 20.0 69.1 66.0
Qwen2.5-VL-7B 50.1 57.3 95.6 23.5 60.7 57.4
16K MiMo-VL-7B-SFT 63.6 71.5 93.2 21.2 68.1 63.5
Qwen2.5-VL-7B 48.7 53.0 91.5 29.1 57.1 55.9
32K MiMo-VL-7B-SFT 63.0 65.6 90.5 21.2 67.2 61.5
Qwen2.5-VL-7B 43.2 47.7 78.5 30.8 57.2 51.5
@@ D.5 More Models for Reasoning Comparison
D.6 Positional Embedding Extrapolation Experiments
@@
According to its technical reports [29], Qwen2.5-VL models are pre-trained with a context length of 32K tokens during the “Long-Context Pre-Training” stage. Meanwhile, its HuggingFace model card shows that we can use YaRN [37], with a scaling factor of 4, to extend its context length to 128K tokens [^7] https://huggingface.co/Qwen/Qwen2.5-VL-3B-Instruct#processing-long-texts. Note that YaRN is used in a zero-shot manner, as there is no continual training for YaRN on Qwen2.5-VL. In Figure˜16, we test the performance of using YaRN. We have two observations: (1) Using YaRN may hurt the performance on shorter input lengths. For example, at 8K tokens, the DocVQA score of Qwen2.5-VL-32B decreases from 67.8% to 63.4%; (2) On average across the entire MMLongBench, YaRN only substantially improves the performance of Qwen2.5-VL-3B (from 21.2 to 30.2 at 128K). However, the SoTA performance from large models (i.e., 32B and 72B) only fluctuates slightly.
@@
Ge et al. [52] proposed a positional embedding extrapolation method called V2PE, where it assigns smaller positional increments to visual tokens than textual tokens. They further applied V2PE to InternVL2-2B and trained the model to enhance its performance on MM-NIAH-Ret (I). In this experiment, we evaluate the V2PE-256K checkpoints [^8] https://huggingface.co/OpenGVLab/V2PE/tree/main/V2PE-256K with the visual increments $\delta\in\{\frac{1}{16},\frac{1}{64},\frac{1}{256}\}$. As shown in Figure˜17, we find that (1) V2PE is very sensitive to different visual increments. For example, when we use $\frac{1}{16}$ and $\frac{1}{64}$, the performance is even worse than InternVL2-2B, which leads to extra hyperparameter tuning; (2) The V2PE (256) shows extremely high performance on NIAH tasks, which can be attributed to the fact that it was trained on MM-NIAH. The sharp performance difference on NIAH tasks versus other categories suggests that V2PE is strongly overfitted to the MM-NIAH dataset.
D.7 Lost in the Middle
@@
Existing works found that text-pure LLMs often struggle to recall needles in the middle of the input sequence, named lost in the middle [57]. On our benchmark, we extend the previous analysis to vision-language tasks with input length up to 128K tokens. We place the needle at six different evenly spaced depths in the context (i.e., $[0,0.2,0.4,0.6,0.8,1.0]$) and evaluate the LCVLMs’ ability to retrieve it. In our study, the needle may be a gold passage, an image, or a key sentence. We show the results in Figure˜18 for ViQuAE, Figure˜19 for VH-Single, Figure˜20 for MM-NIAH-Ret (T), Figure˜21 for MM-NIAH-Ret (I), Figure˜22 for MM-NIAH-Reason (I).
@@
We observe a similar lost-in-the-middle phenomenon in many LCVLMs on long-context vision-language tasks. For example, the InternVL3-14B in Figure˜20 and Ovis2-34B in Figure˜21 both exhibit much better performance when the needle is at depths 0 and $1.0$. Furthermore, as we extend the context to longer lengths (e.g., 128K tokens), we observe cases where the model tends to favor either the very beginning or the very end of the context, but not both simultaneously. For example, as shown in Figure˜18, InternVL3-8B prefers the very beginning of the context (depth 0) at 128K tokens, whereas Qwen2.5-VL-72B favors the very end (depth $1.0$).
D.8 Error Analysis Details
@@
We conducted two error analyses in Section˜4.4. We provide more details of those two analyses in this section. First, for MMLongBench-Doc, we used PyMuPDF  [^9] https://pymupdf.readthedocs.io to extract the plain text from PDF-formatted documents. For ViQuAE, the entity names are already provided in the dataset, since it is constructed based on TriviaQA. The text-pure LLMs we used, corresponding to Qwen2.5-VL models, are Qwen2.5-7B-Instruct and Qwen2.5-32B-Instruct, which are instruction-tuned versions.
D.9 Full Model Evaluation Results
@@
In Figure˜23, we provide the results of all 46 models. We also plot the performance of all 46 models on each dataset in Figures˜24, 25, 26 and 27.
D.10 Idefics2 Performance
@@
The Idefics2-8B and Idefics-8B-C only have a training context window of 2K tokens [49]. We find this leads to very poor long-context generalization. Also, the LLM used in Idefics2 is Mistral-7B-v0.1 [32], whose training length is only 8K tokens. From Figure˜20, we observe that Idefics2 models perform well only when the needle depth is $1.0$ and the context is short (8K or 16K tokens). Additionally, we conduct a sanity check by removing all negative images and retaining only the needle images in Visual Haystack (i.e., one image for single-needle examples and two or three images for multi-needle examples). As shown in Table˜13, we observe that both models achieve performance much higher than a random guess (50%), indicating the correctness of the implementation.
@@
Table 13: Sanity check of Idefics2-8B and Idefics2-8B-C. Here, we use the Visual Haystack dataset. We remove all negative images and only retain needle images (i.e., one image for single-needle examples and two or three images for multi-needle examples).
@@ Appendix E Prompts and Data Examples
We list a few data examples with prompts in Figures˜28, 29, 30, 31, 32 and $3$. For NIAH tasks, we provide examples of both VH and $M$-NIAH, as their input formats are very different.
@@ Appendix F Limitation
[…]
For evaluating summarization, we use a model-based metric (See Section˜A.4) that can provide much better alignment with human judgment than N-gram overlap metrics, such as ROUGE-L. However, we find using GPT-4o to provide the evaluation is expensive, which prevents the long-context community from conducting evaluations with hundreds or even thousands of models. Therefore, it is necessary to find an alternative evaluation method with a lower cost.
@@ Appendix G Broader Impact
The long-context ability of LVLMs has unlocked a large range of applications, including understanding documents with hundreds of pages and reasoning over dozens of web pages automatically. This ability can also help users to summarize a long document or revise a large-scale code repository. Meanwhile, there are a large number of instruction-following scenarios grounded in complex vision-language contexts, such as long-term dialogue with humans or dialogue-based navigation for robots. Looking ahead, our MMLongBench will serve as a standard evaluation for the whole community to benchmark new LCVLMs and to stimulate the development of models with more efficient vision‐language token encodings, more robust position‐extrapolation schemes, and improved OCR, multi‐modal retrieval, and reasoning capabilities.
@@
Figure 30: Example of $M$-NIAH-Ret dataset in NIAH category.